# A Survey on Music Generation from Single-Modal, Cross-Modal, and Multi-Modal Perspectives: Data, Methods, and Challenges

### v1

@@ Abstract.
Multi-modal music generation, using multiple modalities like images, video, and text alongside musical scores and audio as guidance, is an emerging research area with broad applications. This paper reviews this field, categorizing music generation systems from the perspective of modalities. It covers modality representation, multi-modal data alignment, and their utilization to guide music generation. We also discuss current datasets and evaluation methods. Key challenges in this area include effective multi-modal integration, large-scale comprehensive datasets, and systematic evaluation methods. Finally, we provide an outlook on future research directions focusing on multi-modal fusion, alignment, data, and evaluation.
@@
ccs: Applied computing Sound and music computingccs: Computing methodologies Artificial intelligenceccs: General and reference Surveys and overviewsccs: Information systems Multimedia content creation
@@ 2.2. Symbolic Music
Since 1957, researchers have been exploring automated composition (Briot et al., 2020). With the continuous development of symbolic music generation research, numerous encoding methods for symbolic music have been proposed. The event-based symbolic music representation method represents music as a sequence of multiple events according to timestamps, such as note on and note off. The Musical Instrument Digital Interface (MIDI) is a representative of this data type and is widely used in the understanding and generation of symbolic music. However, the MIDI representation may pose difficulties for models learning musical information, such as the necessity for Note on and Note off to appear in pairs. If a model cannot understand this regularity, it cannot correctly generate music. To address the issues with MIDI in model training, Huang and Yang (Huang and Yang, 2020) proposed REMI based on MIDI, resulting in a symbolic music representation more suitable for model learning. For event-based music data, the simplest processing method is to directly treat the event sequence as a 1D sequence. However, this approach often results in excessively long sequences, making it difficult for models to focus on all sequence information. Moreover, a large number of different types of events mixed in the same sequence can make it difficult for models to learn from the data. Therefore, Hsiao et al. (Hsiao et al., 2021) classified music events into note types and metric types, grouping a series of similar events into compound words (CP), which helps models better understand musical information while shortening the sequence length. The OctupleMIDI encoding method (Zeng et al., 2021) further compresses the sequence length compared to CP. Ren et al. (Ren et al., 2020) introduced a representation method suitable for multitrack symbolic music, allowing for the representation of multitrack compositions using a singular symbolic sequence. Roberts et al. (Roberts et al., 2018) introduced a hierarchical decoder on the basis of VAEs for sequential data in MusicVAE, constructing a symbolic music encoding model with good reconstruction and interpolation performance. Additionally, some works represent symbolic music in formats such as piano roll representation (Di et al., 2021; Wang and Yang, 2019) and text (ABC notation) (Wu and Sun, 2023), to fully utilize different model architectures like CNNs and LMs.
@@ 2.3. Text
Natural language text is the primary way humans describe the world. Similarly, a piece of music can also be richly described through text. Advances in NLP technology, particularly the development of LLMs in recent years, have provided abundant solutions and technical foundations for text-to-music task.
@@ 3. Modality-Oriented Music Generation
Existing music generation techniques can derive guidance from multiple modalities, including visual, auditory, and textual inputs, to create corresponding music. It is evident that the content and representation methods of these guidance modalities vary significantly. Therefore, specialized techniques are required for each modality to maximize the utilization of its information and generate high-quality music that meets the desired criteria. This section will categorize and elaborate on the technical details of music generation methods tailored to different guidance modalities.
@@ 3.1. Single-Modal Music Generation
Single-modal music generation leverages existing musical content to produce new compositions within the same modality, encompassing both symbolic-to-symbolic and audio-to-audio generation approaches. This section only discusses part of the representative methods in this field.
@@ 3.1.2. Symbolic-to-Symbolic Generation
While music continuation and inpainting focus on temporal generation, accompaniment generation requires harmonic generation, modeling relationships between simultaneous notes. Zhu et al. (Zhu et al., 2018) presented an end-to-end framework for melody and accompaniment generation. Through multitask learning, the system generates melodies and accompaniment from chord progressions while enabling multi-instrument harmonies. Building on this, the researchers later integrated style discriminators to enable genre-specific generation (Zhu et al., 2020). For long-term accompaniment generation, Ren et al. (Ren et al., 2020) introduced PopMAG, along with MuMIDI, an innovative multi-track MIDI representation. This representation encodes multiple instruments into a unified sequence while explicitly capturing inter-track dependencies. To address sequence length challenges from MuMIDI, the researchers compressed the representations and introduced long-term memory mechanisms inspired by Transformer-XL (Dai et al., 2019). For real-time applications, Wang et al. (Wang et al., 2022) implemented a two-stage generation system. This method distills Transformer (Vaswani et al., 2017) with a CRF (Lafferty et al., 2001) model to maintain high generation quality while keeping low physical latency. Building on this, the researchers later incorporated emotion control through downsampling and semi-supervised learning for smooth affective music adaptation (Wang et al., 2024c).
@@ 3.2. Cross-Modal Music Generation
Manzelli et al. (Manzelli et al., 2018) 2018 LSTM Waveform Mapping WaveNet (Dieleman et al., 2016)
@@ 3.2.1. Score-to-Audio Generation
With advances in deep learning, Manzelli et al. (Manzelli et al., 2018) adapted WaveNet (Dieleman et al., 2016) for score-to-audio synthesis, while Jong Wook Kim et al. enhanced WaveNet with timbre embeddings and FiLM layers for flexible timbre control. However, as a general audio synthesis model, WaveNet lacks specialized design for polyphonic structures of music, complicating the synthesis of harmonically rich audio. To address this, Wang and Yang (Wang and Yang, 2019) introduced a two-stage framework. In the first stage, a U-Net (Ronneberger et al., 2015) architecture learns the mappings between piano roll representations and spectral features. Based on this, a texture refinement network is employed to enhance harmonic details.
@@ 3.3. Multi-Modal Music Generation
Seed-Music (Bai et al., 2024) 2024 Text + Audio + Symbolic → Audio MuLan + VQ-VAE + REMI + Lead sheet Multi-stage conditioning [^3] Including prefix conditioning, cross-attention conditioning, and temporal conditioning. LM + DiT
@@ 3.3.2. Multi-Modal Music Generation Guided by Multiple External Modalities
MuMu-LLaMA (Liu et al., 2024b) is a comprehensive vision-music understanding and generation model capable of generating music guided by images, videos and text. It employs ViT (Dosovitskiy et al., 2021) for image understanding and ViViT (Arnab et al., 2021) for video understanding. LLaMA 2 (Touvron et al., 2023) acts as a bridge for multi-modal understanding, trained to process and integrate visual, textual and audio information. During music generation, visual encodings from the understanding modules are fed into LLaMA 2, and the final layer of LLaMA 2 is passed to an output projector. This projector converts the outputs of LLaMA 2 into a music description embedding vector, which serves as a conditioning signal for pre-trained music generation models like AudioLDM (Liu et al., 2023, 2024c) and MusicGen (Copet et al., 2024). Finally, the pre-trained model synthesizes the corresponding music based on the embedding vector.
Seed-Music (Bai et al., 2024) is a unified multi-modal music generation system integrating text, audio, and symbolic modalities. For text encoding, the system encodes textual descriptions via MuLan and represents lyrics as phonemes. For audio encoding, it uses LM-based audio tokenizers. For symbolic encoding, it employs REMI (Huang and Yang, 2020) as notation and uses lead sheets as tokens. Seed-Music comprises three complementary pipelines working together. The LM-based pipeline uses Transformer to generate audio tokens which are then rendered by a DiT model (Peebles and Xie, 2023). The symbolic pipeline generates and renders lead sheet tokens instead, enhancing interpretability and editing capabilities. The vocoder pipeline directly generates audio from input conditions using DiT. By combining these pipelines, Seed-Music integrates text, audio, and symbolic modalities while enabling user editing. The study also compares LM-based and diffusion-based pipelines, finding LM-based approaches more suitable for user interaction and future integration with other multi-modal models.
@@ 4.3. Visual-Music Datasets
Beyond performance and sport videos, the real world contains even richer and more complex videos. Therefore, the field of video-to-music generation requires more generalized video-music datasets. In reality, platforms like YouTube and TikTok  [^11] https://www.tiktok.com host large amounts of video data. However, many of these data are of low quality, presenting a considerable challenge for data cleaning. A potential solution to address this issue is to filter for Music Videos (MVs), which typically exhibit strong video-music consistency and generally maintain high quality. Hong et al. (Hong et al., 2018) filtered MVs from the YouTube-8M dataset (Abu-El-Haija et al., 2016), resulting in HIMV-200K. HIMV-200K contains 200,500 video-music pairs, including official MVs, parody MVs, and user-generated MVs, offering strong diversity and complexity. However, HIMV-200K lacks fine-grained annotations for musical features and still includes many low-quality data points. To overcome these issues, Zhuo et al. (Zhuo et al., 2023) constructed the SymMV dataset, which contains 1,140 video-audio pairs totaling 76.5 hours. The researchers first retrieved professional pianist performance videos from YouTube and then searched for the corresponding official MVs based on the metadata. The piano performance audio is then transcribed into MIDI using a piano transcription model (Hawthorne et al., 2018), enabling fine-grained musical annotations.
@@ 4.5. Solutions for Multi-Modal Data Scarcity
Employing pre-training or multi-stage methods can enable models to learn from single-modal data, reducing the need for highly aligned multi-modal data. SongMASS (Sheng et al., 2021) introduces pre-training methods to allow models to learn from unpaired single-modal data, using only a small amount of paired data to train the model’s cross-modal transfer capabilities. TeleMelody (Ju et al., 2022) introduces intermediate representations for two modalities, splitting cross-modal tasks so that model training no longer relies on paired data.
Rule-based modal mapping methods can also enable models to acquire cross-modal capabilities. CMT (Di et al., 2021) introduces a rule-based mapping relationship between video rhythm and music. Following CMT, XMusic (Tian et al., 2025) utilizes mapping algorithms and pre-training understanding models to map various modalities to musical features. The two models are both trained on single-modal datasets, but they can be applied to cross-modal and multi-modal music generation through rule-based mapping methods.
@@ 5. Evaluation
Currently, there exist a number of multi-modal music generation algorithms, making the evaluation of these methods a significant concern. However, music is an art that is constrained by objective factors such as acoustics and music theory, while also being influenced by subjective factors like personal experiences, expertise, aesthetics, and emotions of the listeners. Thus, assessing the quality of music is not straightforward. It requires both an objective evaluation of the music’s rationality based on theoretical grounds and a subjective evaluation from human listeners. Moreover, multi-modal music generation emphasizes the use of information from the guidance modalities to produce music that meets specific requirements. Consequently, the evaluation of multi-modal music generation extends beyond the assessment of inherent musical quality. A crucial dimension is the evaluation of cross-modal consistency, which quantifies the fidelity with which the generated music reflects the semantic and structural information present in the input modalities. These aspects pose challenges to evaluation methods, necessitating the design of robust evaluation systems to assess multi-modal music generation approaches. Although there is no unified evaluation system in the field of multi-modal music generation, researchers are striving to propose more reasonable and efficient evaluation metrics. This section will introduce the widely used subjective and objective metrics for multi-modal music generation.
@@ 5.1.3. Consistency
Evaluation Based on Contrastive Embedding Models. CLAP Score (Elizalde et al., 2023) assesses text-audio alignment by calculating the cosine similarity between the generated music and the textual description in the joint embedding space of CLAP. A higher score indicates a better match between the generated music and the textual description. Similarly, MuLan Cycle Consistency (MCC) (Agostinelli et al., 2023) measures the semantic consistency by computing the cosine similarity in the MuLan embedding space (Huang et al., 2022a). For evaluating the semantic consistency between generated music and input images, Chowdhury et al. (Chowdhury et al., 2024) proposed a new metric, Image Music Similarity Metric (IMSM). IMSM combines the text-image embedding model CLIP (Radford et al., 2021) and the text-audio embedding model CLAP (Elizalde et al., 2023), using text as a bridge to connect image and audio semantics. Similarly, Liu et al. (Liu et al., 2024b) introduced ImageBind Ranking (IB Rank), using the ImageBind model (Girdhar et al., 2023) to measure the relevance between generated music and the input images or videos. ImageBind is a multi-modal alignment model capable of mapping different modalities into the same embedding space, enabling cross-modal semantic alignment. In IB Rank, the embeddings of the input and the generated music are first extracted using ImageBind, and then the cosine similarity between these embeddings is calculated. However, it is important to note that such evaluation methods relying on joint embedding models depend on the quality of the pre-trained models. Additionally, although these methods can directly measure semantic alignment, they cannot reflect other characteristics of the music.

### v2

# A Survey on Music Generation from Single-Modal, Cross-Modal and Multi-Modal Perspectives

@@ Abstract.
Multi-modal music generation, using multiple modalities like text, images, and video alongside musical scores and audio as guidance, is an emerging research area with broad applications. This paper reviews this field, categorizing music generation systems from the perspective of modalities. The review covers modality representation, multi-modal data alignment, and their utilization to guide music generation. Current datasets and evaluation methods are also discussed. Key challenges in this area include effective multi-modal integration, large-scale comprehensive datasets, and systematic evaluation methods. Finally, an outlook on future research directions is provided, focusing on creativity, efficiency, multi-modal alignment, and evaluation.
@@
This work is supported by the National Natural Science Foundation of China (No.62272409)
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2025/04ccs: Applied computing Sound and music computingccs: Computing methodologies Artificial intelligenceccs: General and reference Surveys and overviewsccs: Information systems Multimedia content creation
@@ 2.2. Symbolic Music
Since 1957, researchers have been exploring automated composition (Briot et al., 2020). With the continuous development of symbolic music generation research, numerous encoding methods for symbolic music have been proposed. The event-based symbolic music representation method represents music as a sequence of multiple events according to timestamps, such as note on and note off. The Musical Instrument Digital Interface (MIDI)  [^2] https://midi.org/ is a representative of this data type and is widely used in the understanding and generation of symbolic music. However, the MIDI representation may pose difficulties for models learning musical information, such as the necessity for Note on and Note off to appear in pairs. If a model cannot understand this regularity, it cannot correctly generate music. To address the issues with MIDI in model training, Huang and Yang (Huang and Yang, 2020) proposed REMI based on MIDI, resulting in a symbolic music representation more suitable for model learning. For event-based music data, the simplest processing method is to directly treat the event sequence as a 1D sequence. However, this approach often results in excessively long sequences, making it difficult for models to focus on all sequence information. Moreover, a large number of different types of events mixed in the same sequence can make it difficult for models to learn from the data. Therefore, Hsiao et al. (Hsiao et al., 2021) classified music events into note types and metric types, grouping a series of similar events into compound words (CP), which helps models better understand musical information while shortening the sequence length. The OctupleMIDI encoding method (Zeng et al., 2021) further compresses the sequence length compared to CP. Ren et al. (Ren et al., 2020) introduced a representation method suitable for multitrack symbolic music, allowing for the representation of multitrack compositions using a singular symbolic sequence. Roberts et al. (Roberts et al., 2018) introduced a hierarchical decoder on the basis of VAEs for sequential data and developed MusicVAE, constructing a symbolic music encoding model with good reconstruction and interpolation performance. Additionally, some works represent symbolic music in formats such as piano roll representation (Di et al., 2021; Wang and Yang, 2019) and text (ABC notation) (Wu and Sun, 2023), to fully utilize different model architectures like CNNs and LMs.
@@ 2.3. Text
Natural language text is the primary way humans describe the world. Similarly, a piece of music can also be richly described through text. Advances in NLP technology, particularly the development of LLMs in recent years, have provided abundant solutions and technical foundations for text-to-music generation.
@@ 3. Modality-Oriented Music Generation
Existing music generation techniques can derive guidance from multiple modalities, including auditory, textual, and visual inputs, to create corresponding music. It is evident that the content and representation methods of these guidance modalities vary significantly. Therefore, specialized techniques are required for each modality to maximize the utilization of its information and generate high-quality music that meets the desired criteria. This section will categorize and elaborate on the technical details of music generation methods tailored to different guidance modalities.
@@ 3.1. Single-Modal Music Generation
Single-modal music generation leverages existing musical content to produce new compositions within the same modality, encompassing both symbolic-to-symbolic and audio-to-audio generation approaches. This section discusses part of the representative methods in this field.
@@ 3.1.2. Symbolic-to-Symbolic Generation
While music continuation and inpainting focus on temporal generation, accompaniment generation requires harmonic generation, modeling relationships between simultaneous notes. Zhu et al. (Zhu et al., 2018) presented an end-to-end framework for melody and accompaniment generation. Through multitask learning, the system generates melodies and accompaniment while enabling multi-instrument harmonies. Building on this, the researchers integrated style discriminators to enable genre-specific generation (Zhu et al., 2020). For long-term accompaniment generation, Ren et al. (Ren et al., 2020) introduced PopMAG, along with MuMIDI, an innovative multi-track MIDI representation. This representation encodes multiple instruments into a unified sequence while explicitly capturing inter-track dependencies. To address sequence length challenges from MuMIDI, the researchers compressed the representations and introduced long-term memory mechanisms inspired by Transformer-XL (Dai et al., 2019). For real-time applications, Wang et al. (Wang et al., 2022) implemented a two-stage generation system. This method distills Transformer (Vaswani et al., 2017) with a CRF (Lafferty et al., 2001) model to maintain high generation quality while keeping low physical latency. The researchers later extended the work towards multi-modal music generation, incorporating emotion control through downsampling and semi-supervised learning for smooth affective music adaptation (Wang et al., 2024c).
@@ 3.2. Cross-Modal Music Generation
Manzelli et al. (Manzelli et al., 2018) 2018 MIDI + LSTM Waveform Mapping WaveNet (Dieleman et al., 2016)
Mel2Mel (Kim et al., 2019) 2019 Onsets and Frames (Hawthorne et al., 2018) Mel spectrogram Mapping Bi-LSTM + WaveNet (Dieleman et al., 2016)
@@
MusTango (Melechovsky et al., 2024a) 2024 FLAN-T5 + Embedding VAE Cross-attention LDM
MeLoDy (Lam et al., 2024) 2024 MuLan Wav2Vec2-Conformer (Baevski et al., 2020; Gulati et al., 2020) + VAE-GAN Joint embedding LLaMA + LDM
@@ 3.2.1. Score-to-Audio Generation
With advances in deep learning, Manzelli et al. (Manzelli et al., 2018) adapted WaveNet (Dieleman et al., 2016) for score-to-audio synthesis, while Kim et al. (Kim et al., 2019) enhanced WaveNet with timbre embeddings and FiLM layers for flexible timbre control. However, as a general audio synthesis model, WaveNet lacks specialized design for polyphonic structures of music, complicating the synthesis of harmonically rich audio. To address this, Wang and Yang (Wang and Yang, 2019) introduced a two-stage framework. In the first stage, a U-Net (Ronneberger et al., 2015) architecture learns the mappings between piano roll representations and spectral features. Based on this, a texture refinement network is employed to enhance harmonic details.
@@ 3.3. Multi-Modal Music Generation
Seed-Music (Bai et al., 2024) 2024 Text + Audio + Symbolic → Audio MuLan + VQ-VAE + REMI + Lead sheet Multi-stage conditioning* LM + DiT
@@
    Including prefix conditioning, cross-attention conditioning, and temporal conditioning.
@@ 3.3.1. Multi-Modal Music Generation Guided by a Single External Modality
However, such methods solely rely on a single external modality to guide the transformation of internal modalities, which is insufficient for exploring multi-modal interactions.
@@ 3.3.2. Multi-Modal Music Generation Guided by Multiple External Modalities
Seed-Music (Bai et al., 2024) is a unified multi-modal music generation system integrating text, audio, and symbolic modalities. For text encoding, the system encodes textual descriptions via MuLan and represents lyrics as phonemes. For audio encoding, it uses LM-based audio tokenizers. For symbolic encoding, it employs REMI (Huang and Yang, 2020) as notation and uses lead sheets as tokens. Seed-Music comprises three complementary pipelines working together. The LM-based pipeline uses Transformer to generate audio tokens which are then rendered by a DiT model (Peebles and Xie, 2023). The symbolic pipeline generates and renders lead sheet tokens instead, enhancing interpretability and editing capabilities. The vocoder pipeline directly generates audio from input conditions using DiT. By combining these pipelines, Seed-Music integrates text, audio, and symbolic modalities while enabling user editing. The study also compares LM-based and diffusion-based pipelines, finding LM-based approaches more suitable for user interaction and future integration with other multi-modal models. Although Seed-Music integrates text, symbolic music, and audio modalities, it lacks guidance from the visual modality.
@@
While these works jointly leverage visual and textual modalities to guide music generation, they lack a unified framework to integrate additional modalities. MelFusion solely employs images to provide visual information, without incorporating video. Although Diff-BGM integrates both video and textual information, the textual modality is merely used to express video semantics, resulting in insufficient exploration of text. Furthermore, it introduces control signals from the two modalities via feature selection. Only one modality guides music generation at any given time, leading to inadequate fusion between the two. V2Meow incorporates limited textual information, offering minimal guidance for music generation. The unified integration of multiple modalities remains a critical trend in multi-modal research.
MuMu-LLaMA (Liu et al., 2024b) is a comprehensive visual-music understanding and generation model capable of generating music guided by images, videos and text. It employs ViT (Dosovitskiy et al., 2021) for image understanding and ViViT (Arnab et al., 2021) for video understanding. LLaMA 2 (Touvron et al., 2023) acts as a bridge for multi-modal understanding, trained to process and integrate visual, textual and audio information. During music generation, visual encodings from the understanding modules are fed into LLaMA 2, and the final layer of LLaMA 2 is passed to an output projector. This projector converts the outputs of LLaMA 2 into a music description embedding vector, which serves as a conditioning signal for pre-trained music generation models like AudioLDM (Liu et al., 2023, 2024c) and MusicGen (Copet et al., 2024). Finally, the pre-trained model synthesizes the corresponding music based on the embedding vector.
@@ 4.3. Visual-Music Datasets
Beyond performance and sports, the real world contains even richer videos. Video-to-music generation requires more generalized video-music datasets. In reality, platforms like YouTube and TikTok  [^11] https://www.tiktok.com host large amounts of video data. However, many of these data are of low quality, presenting a considerable challenge for data cleaning. A potential solution to address this issue is to filter for Music Videos (MVs), which typically exhibit strong video-music consistency and generally maintain high quality. Hong et al. (Hong et al., 2018) filtered MVs from the YouTube-8M dataset (Abu-El-Haija et al., 2016), resulting in HIMV-200K. HIMV-200K contains 200,500 video-music pairs, including official MVs, parody MVs, and user-generated MVs, offering strong diversity and complexity. However, HIMV-200K lacks fine-grained annotations for musical features and still includes many low-quality data points. To overcome these issues, Zhuo et al. (Zhuo et al., 2023) constructed the SymMV dataset, which contains 1,140 video-audio pairs totaling 76.5 hours. The researchers first retrieved professional pianist performance videos from YouTube and then searched for the corresponding official MVs based on the metadata. The piano audio clips were transcribed into MIDI files, enabling fine-grained musical annotations.
@@ 4.5. Solutions for Multi-Modal Data Scarcity
Enhancing model architectures or training methods can enable models to learn from single-modal data, reducing the need for highly aligned multi-modal data. SongMASS (Sheng et al., 2021) introduces pre-training methods to allow models to learn from unpaired single-modal data, using only a small amount of paired data to train the model’s cross-modal transfer capabilities. TeleMelody (Ju et al., 2022) introduces intermediate representations for two modalities, splitting cross-modal tasks so that model training no longer relies on paired data. CMT (Di et al., 2021) introduces a rule-based mapping relationship between video rhythm and music. Following CMT, XMusic (Tian et al., 2025) utilizes mapping algorithms and pre-training understanding models to map various modalities to musical features. The two models are both trained on single-modal datasets, but they can be applied to cross-modal and multi-modal music generation through rule-based mapping methods.
@@ 5. Evaluation
Currently, there exist a number of multi-modal music generation algorithms, making the evaluation of these methods a significant concern. However, music is an art that is constrained by objective factors such as acoustics and music theory, while also being influenced by subjective factors like personal experiences, expertise, and aesthetics. Thus, assessing the quality of music is not straightforward. It requires both an objective evaluation based on theoretical grounds and a subjective evaluation from human listeners. Moreover, multi-modal music generation emphasizes the use of information from the guidance modalities to produce music that meets specific requirements. Consequently, the evaluation of multi-modal music generation extends beyond the assessment of inherent musical quality. A crucial dimension is the evaluation of cross-modal consistency, which quantifies the fidelity with which the generated music reflects the input modalities. These aspects pose challenges to evaluation methods, necessitating the design of robust evaluation systems to assess multi-modal music generation approaches. Although there is no unified evaluation system in the field of multi-modal music generation, researchers are striving to propose more reasonable and efficient evaluation metrics. This section will introduce the widely used subjective and objective metrics for multi-modal music generation.
@@ 5.1.3. Consistency
Evaluation Based on Contrastive Embedding Models. CLAP Score (Elizalde et al., 2023) assesses text-audio alignment by calculating the cosine similarity between the generated music and the textual description in the joint embedding space of CLAP. A higher score indicates a better match between the generated music and the textual description. Similarly, MuLan Cycle Consistency (MCC) (Agostinelli et al., 2023) measures the semantic consistency by computing the cosine similarity in the MuLan embedding space (Huang et al., 2022a). For evaluating the semantic consistency between generated music and input images, Chowdhury et al. (Chowdhury et al., 2024) proposed a new metric, Image Music Similarity Metric (IMSM). IMSM combines the text-image embedding model CLIP (Radford et al., 2021) and the text-audio embedding model CLAP (Elizalde et al., 2023), using text as a bridge to connect image and audio semantics. Similarly, Liu et al. (Liu et al., 2024b) introduced ImageBind Ranking (IB Rank), using the ImageBind model (Girdhar et al., 2023) to measure the relevance between generated music and the input images or videos. ImageBind is a multi-modal alignment model capable of mapping different modalities into the same embedding space, enabling cross-modal semantic alignment. In IB Rank, the embeddings of the input and the generated music are first extracted using ImageBind, and then the cosine similarity between these embeddings is calculated. However, it is important to note that such evaluation methods relying on joint embedding models depend on the quality of the pre-trained models. Furthermore, although these methods can directly measure semantic alignment, they cannot reflect other characteristics of the music.